# Detecting Adversarial Data via Perturbation Forgery

### v2

@@ Introduction
Table 6: Time cost of inference (second) on ImageNet100
Detector LID LiBRe EPSAD SPAD ours
Time (second) 1.80 2.56 396.81 4.56 4.85
Detecting Face Adversarial Examples
Adversarial attack on the face recognition model is an important branch of adversarial example research. Therefore, the detection ability of adversarial face examples is an important indicator of the performance of the detector. We perform various adversarial face attacks and compare our approach with state-of-the-art methods from adversarial face detection and face forgery detection. As shown in Table 4, our method shows excellent performance against a variety of adversarial attacks targeting faces and surpasses many existing adversarial face detection techniques. It further indicates the ability of the sparse mask to capture the essential characteristics of face adversarial examples and the strong generalization capability of our method.
Ablation Study and Analysis
To further evaluate our method, we conduct experiments to show the effect of the sparse mask and compare the time cost to baseline methods. We also analyze the functionality of Perturbation Forgery through visualizations. The impact of hyper-parameters is reported in the supplementary materials.
Impact of each mask
@@
As argued, the sparse mask is proposed to assist the detector in identifying local adversarial noises present in generative-based adversarial data. As shown in Table 5, the absence of the saliency detection model $f_{s}$ or the GradCAM model $f_{c}$ has little effect on the overall system. This is because the regions masked by $f_{s}$ and $f_{c}$ are somewhat complementary. However, without the Sobel operation to make the mask sparse, the detector fails entirely in detecting attacks, except for CDA, because the noise in generative-based adversarial data tends to be imbalanced and localized, whereas the noise in CDA is more global. Without the sparsification provided by the Sobel operation, the generated pseudo-noise becomes too dense, preventing the detector from effectively focusing on local noise.
Time Cost of Inference
We conducted experiments to calculate the inference time cost for 100 samples on ImageNet100. As shown in Table 6, the time cost of our model is slightly higher than that of LID, LiBRe, and SPAD, and it only takes 0.0485 seconds to process one image. For actual use, a minor increase in time is perfectly acceptable in exchange for a significant increase in detection performance.
Figure 3: 2D T-SNE visualizations. (a) ImageNet100 flattened noises of adversarial data and Perturbation Forgery. (b) ImageNet100 features extracted by the model trained with Perturbation Forgery.
Functionality of Perturbation Forgery
@@
To validate the functionality of our method, we extract features and flattened noises from adversarial data and Perturbation Forgery and visualize them using 2D t-SNE projection. All adversarial data come from unseen attacks. As shown in Figure 3(a), by continuously perturbing the adversarial noise distribution of a commonly used attack, Perturbation Forgery forms an open covering of noises from unseen attacks. Individual samples float outside of the collective because they are outliers from the noise distribution. As shown in Figure 3(b), by learning to separate natural data from those generated by Perturbation Forgery, the trained detector can effectively distinguish between natural and adversarial data.
Conclusion
@@
In this paper, we explore and define the proximity relationship between adversarial noise distributions by modeling these distributions as multivariate Gaussian. Based on our assumptions and definitions, we establish a perturbation method to craft proximal distributions from a given distribution and demonstrate the existence of an open covering of adversarial noise distributions. After that, we propose the core idea of obtaining a detector with strong generalization abilities against all types of adversarial attacks by learning to distinguish the open covering from the distribution of natural data. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, agnostic to any specific models. The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging. We must rely on continuous experimentation to determine these parameters. However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
Appendix
Appendix A Proofs
Theorem Proofs
Proof of Theorem 1.
Proof.
Denoting the distribution of natural sample $\mathbf{x}$ as $P$, we use the 1-Wasserstein distance as the metric $d(\cdot,\cdot)$, to prove the theorem.
1) We denote the independent adv-noise distribution of $\eta_{i}$ as $Q_{i}\in\mathcal{P}_{a}$ where $\eta_{i}=\hat{\mathbf{x}}_{i}-\mathbf{x}$, and a zero-distribution as $P$ where its mean and variance are zero vectors. Then the 1-Wasserstein distance between $\eta_{i}$ and $\eta_{j}$ can be bound by its metric property and dual form:
$W_{1}\left(Q_{i},Q_{j}\right)\leq$ $W_{1}(Q_{i},P)+W_{1}(Q_{j},P)$
$\leq$ $\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{i}\sim Q_{i}}[g(\boldsymbol{\eta}_{i})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
$+\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{j}\sim Q_{j}}[g(\boldsymbol{\eta}_{j})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
$\leq$ $2\epsilon,$
where $\|g\|_{L}$ is the Lipschitz constant of function $g$ w.r.t. the norm induced by the metric here.
2) Consider two random noise $\boldsymbol{\eta}_{i}\sim\mathcal{N}_{i}(\mu_{i},\sigma_{i}^{2}\mathbf{I}_{d})$ and $\boldsymbol{\eta}_{j}\sim\mathcal{N}_{j}(\mu_{j},\sigma_{j}^{2}\mathbf{I}_{d})$. While it’s mean and variance satisfied that $\|\mu_{i}-\mu_{j}\|\leq\epsilon$ and $|\sigma_{i}-\sigma_{j}|\leq\epsilon$. then the 2-Wasserstein distance between $\eta_{1}$ and $\eta_{2}$ is
$W_{2}\left(\boldsymbol{\eta}_{i},\boldsymbol{\eta}_{j}\right)^{2}=$ $\left\|\mu_{i}-\mu_{j}\right\|^{2}+\operatorname{trace}(\sigma_{i}^{2}\mathbf{I}_{d}+\sigma_{j}^{2}\mathbf{I}_{d}$
$-2\left((\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\sigma_{i}^{2}\mathbf{I}_{d}(\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\right)^{1/2})$
$\leq$ $\epsilon_{\mu}^{2}+d\cdot\epsilon_{\sigma}^{2},$
where $(\sigma_{i}^{2}\mathbf{I}_{d})^{1/2}$ denotes the principal square root of $\sigma_{i}^{2}\mathbf{I}_{d}$. ∎
Corollary 1.
1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$.
2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$.
3) Based on the properties of the complete metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists a finite open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
Proof.
1) Consider 1-Wasserstein distance as metric $d$, for $\forall Q_{i},Q_{j}\in\mathcal{P}_{\varepsilon}$, we have
$W_{1}\left(Q_{i}\right)\geq 0.\quad W_{1}\left(Q_{i},Q_{j}\right)=0\ if\ and\ only\ if\ Q_{i}=Q_{j}.$ (13)
$W_{1}\left(Q_{i},Q_{j}\right)=W_{1}\left(Q_{j},Q_{i}\right).$ (14)
$W_{1}\left(Q_{i},Q_{j}\right)\leq W_{1}(Q_{i},P)+W_{1}(Q_{j},P).$ (15)
Therefore, $(\mathcal{P}_{\varepsilon},d)$ is a metric space.
2) It is obvious according to the definition of $\mathcal{P}_{\varepsilon}$.
3) For any subset $\mathcal{P}_{i}=\{p_{i1},p_{i2},\dots,p_{in}\}\subset\mathcal{P}_{\varepsilon}$, we can construct several spherical subsets $\{\mathcal{P}_{ij}\}$ with $p_{i1}$ as the center and $|\varepsilon-d(p_{i1},p(\boldsymbol{\eta}))|$ as the radius. $\{\mathcal{P}_{ij}\}$ is a open covering of $\mathcal{P}_{i}$.
Appendix B Experimental Details
Dataset
CIFAR10
CIFAR10 consists of a training set of 50,000 images and a test set of 10,000 images, each with a resolution of 32x32 pixels. These images are divided into 10 different classes. The classes represent everyday objects such as airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
ImageNet100
ImageNet-100 is a subset of ImageNet (Deng et al. 2009), comprising 100 classes selected from the original ImageNet dataset, each containing a substantial number of high-resolution images, with 1,300 images per class.
Makeup
Makeup comprises images of faces with and without makeup. These images are collected from various sources to ensure a diverse representation of facial features, makeup styles, and skin tones. There are 333 before-makeup images and 302 after-makeup images.
Table 7: Detection AUC(%) against various gradient-based attacks on CIFAR-10 ($\epsilon=4/255$).
Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
SPAD (Wang et al. 2023a) 99.51 99.38 99.35 99.52 99.66 99.85 99.47 99.63 99.74
EPSAD (Zhang et al. 2023) 99.99 99.98 99.98 99.99 99.99 99.99 99.99 99.99 99.99
ours 99.65 99.65 99.68 99.47 99.80 99.80 99.76 99.64 99.88
Figure 4: The influence of parameters on the detection performance.
Figure 5: The influence of attack intensity on the detection performance.
CelebA-HQ
CelebA-HQ dataset is a high-quality version of the original CelebA dataset.CelebA-HQ consists of 30,000 high-resolution images of celebrity faces, derived from the CelebA dataset through a progressive GAN-based upsampling and quality enhancement process. Each image in CelebA-HQ is 1024x1024 pixels.
LFW
LFW contains 13,233 labeled images of 5,749 different individuals collected from the internet, each image is labeled with the name of the person pictured. LFW images vary widely in terms of lighting, facial expression, pose, and background, closely reflecting real-world conditions. The images are provided in a resolution of 250x250 pixels, and the faces are roughly aligned based on the eye coordinates.
Adversarial Attacks
Gradient-based attacks
We use Torchattacks (Kim 2020) to implement adversarial attacks. The victim models used for adversarial data generation are ResNet-50 (He et al. 2016) for ImageNet, WideReseNet-28 for CIFAR10 (Krizhevsky, Hinton et al. 2009), and ArcFace (Deng et al. 2019) for face attacks.
Generative-model-based attacks
For CDA, we use the pre-trained generator which has been trained on ImageNet with a relativistic loss against ResNet152. For TTP, we use the pre-trained-generators trained against ResNet50 for 8 target labels, they are: Grey Owl(24),Goose(99),French Bulldog(245),Hippopotamus(344),Cannon(471),Fire Engine(555),Parachute(701), Snowmobile(802). We generate 1000 adversarial examples for each label and randomly select 125 adversarial examples for each label. For M3D, its settings are consistent with TTP. For Diff-PGD, we use the global attack to craft adversarial examples, target classifier is ResNet50, the diffusion model accelerator is ddim50, the reverse step in SDEdit is 3, and the iteration number of PGD is 10, the step size is 2. For Diff-attack, the target classifier is ResNet50, DDIM sample steps are set as 20, and iterations to optimize the adversarial image are set as 30. For the methods mentioned above, we all use the authors’ implementation to craft adversarial examples.
Face attacks
For face attacks, 1 gradient-based attack TIPIM (Yang et al. 2021), two GAN-based attacks including AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022), and 3 physical attacks including Adv-Sticker (Tong et al. 2021), Adv-Glasses (Sharif et al. 2016), and Adv-Mask (Tong et al. 2021) produce face adversarial samples by attacking the victim face model ArcFace (Deng et al. 2019).
Baselines
Adversarial detection
For LiBRe [^1] https://github.com/thudzj/ScalableBDL/tree/efficient/exps, we trained a ResNet-50 on ImageNet100 and finetuned it using the authors’ implementation, the parameters were kept consistent with theirs. For EPSAD [^2] https://github.com/ZSHsh98/EPS-AD, we use the pre-trained models provided by the authors and follow the instructions by the authors to get the detection results. Due to the time-consuming nature of computing EPS, we evaluated 1,000 images for each attack method on ImageNet-100.
Synthetic image detection
Since existing adversarial example detection methods do not address adversarial examples based on generative models, we compared our method with various synthetic image detection methods to effectively validate its effectiveness. We conducted detection experiments on adversarial examples based on generative models using the following methods: CNN-Detection(Wang et al. 2020) [^3] https://github.com/peterwang512/CNNDetection, LGrad(Tan et al. 2023) [^4] https://github.com/peterwang512/CNNDetection, Universal-Detector(Ojha, Li, and Lee 2023) [^5] https://github.com/Yuheng-Li/UniversalFakeDetect?tab=readme-ov-file#weights, and DIRE(Wang et al. 2023b) [^6] https://github.com/ZhendongWang6/DIRE. We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
Implementation Details
Experiments are implemented using 2 Nvidia Deforce RTX 3090 GPUs. For training detectors, we use an Adam optimizer with a learning rate of 2e-4, momentum of 0.9, and weight decay of 5e-6.
Data pre-processing
We resize images to $32\times 32$ for CIFAR10, $256\times 256$ for ImageNet100, Makeup and CelebA-HQ, and $112\times 112$ for LFW. For ImageNet100, after resizing, we center-crop the images to $224\times 224$. After that, all of the images are resized to $256\times 256$ followed by random horizontal flip and normalization.
Noise distribution
As related in the main paper, we flatten the generated noise to a vector and estimate its distribution. However, for a noise with a shape of $256\times 256\times 3$, the length of the corresponding flattened noise is $196,608$, making it impossible to run the algorithm on a single GPU. Therefore, we fix the size of noise to $32\times 32\times 3$, and concatenate several sampled pseudo noises to fit the larger image.
Figure 6: Natural images and corresponding pseudo-adversarial images generated by Perturbation Forgery.
Experiment on CIFAR-10
We conduct experiments to detect gradient-based attacks on CIFAR10. As reported in Table 7, our method demonstrates effective detection (achieving an AUC score over 0.99) across a variety of attack methods, validating our assumptions of the noise distribution proximity.
Impact of Parameters
We select typical attacks and present curves in Figure 4 to depict the impact of parameters on the detection performance. The parameters $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ have little effect on detecting PGD, while $\gamma_{c}$ and $\gamma_{l}$ have significant effects on detecting M3D and Diff-PGD. This might be because, regardless of whether $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ are large or small, the accuracy of forming the open coverage remains unaffected as long as the data is sufficient. However, too small thresholds lead to adding pseudo-noise to insensitive and low-frequency regions, while too large thresholds result in extremely sparse pseudo-noise that is far from the adversarial noise distribution.
To explore the impact of perturbation intensity of attacks, we select typical attacks and present the curves in Figure 5, where $\epsilon_{\mathrm{a}}$ denotes the intensity of unseen attacks and $\epsilon_{\mathrm{i}}$ denotes the intensity of the initial attack for Perturbation Forgery. We can see that the detector trained with a minor $\epsilon_{\mathrm{i}}$ is able to detect adversarial data generated with a larger $\epsilon_{\mathrm{a}}$. When $\epsilon_{\mathrm{i}}$ is larger, adversarial data generated with a minor $\epsilon_{\mathrm{a}}$ is not easy to detect. This might be because adversarial noises generated with a higher $\epsilon_{\mathrm{a}}$ tend to be more obvious to detect. When $\epsilon_{\mathrm{a}}$ is too small than $\epsilon_{\mathrm{i}}$, the distance between the initial attack noise and the unseen attack noise is too far away, reducing the detection performance.
Figure 7: 2D T-SNE visualizations. (a) CIFAR-10 flattened noises of adversarial data and Perturbation Forgery. (b) CIFAR-10 features extracted by the model trained with Perturbation Forgery.
Appendix C Limitations
The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging. We must rely on continuous experimentation to determine these parameters. Specifically, we initiate parameters of uniform distribution as $\epsilon_{\mu}=1$ and $\epsilon_{\sigma}=0.001$, and adjust them using a small step factor. Finally, we find the experiment results when $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$ are close to optimal. However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
Appendix D More Visualization Results
Visualization on CIFAR-10
We extract features and flattened noises from adversarial data and Perturbation Forgery from CIFAR-10 and visualize them using 2D t-SNE projection in Figure 7.
Visualization Examples
Some examples of Perturbation Forgery at $\epsilon=4/255$ are shown in Figure 6.

### v4

@@ 5.4 Detecting Face Adversarial Examples
5.5 Ablation Study and Analysis
To further evaluate our method, we report an ablation study on the sparse mask, white-box attack detection results, the impact of the initial attack and the data size for distribution estimation, and the functionality of Perturbation Forgery through visualizations. The impact of hyper-parameters and other analyses are reported in the supplementary materials.
5.5.1 Impact of each mask
@@
As argued, the sparse mask is proposed to assist the detector in identifying local adversarial noises in generative-based adversarial data. As shown in Table 5, the absence of the saliency detection model $f_{s}$ or the GradCAM model $f_{c}$ has little effect on the overall system. This is because the regions masked by $f_{s}$ and $f_{c}$ are somewhat complementary. However, without the Sobel operation to sparsify the mask, the detector fails to detect most attacks, except for CDA. This is because generative-based adversarial noise tends to be imbalanced and localized, while CDA’s noise is more global. The generated pseudo-adversarial noise without Sobel becomes too dense, preventing the detector from effectively focusing on localized noise.
5.5.2 Functionality Visualization
@@
To validate the functionality of our method, we extract features and flattened noise from adversarial data and Perturbation Forgery-generated data, then visualize them using 2D t-SNE projections. All adversarial data originate from unseen attacks. As shown in Figure 3(a), Perturbation Forgery continuously perturbs the adv-noise distribution of the initial attack, thereby forming an open covering for noises from unseen attacks. Individual samples appear as outliers, floating outside the collective noise distribution. As shown in Figure 3(b), by learning to separate natural data from those generated by Perturbation Forgery, the trained detector can effectively distinguish between natural and adversarial data.
Figure 3: 2D T-SNE visualizations. (a) ImageNet100 flattened noises of adversarial data and Perturbation Forgery-generated data. (b) ImageNet100 features extracted by the backbone model trained with Perturbation Forgery.
Table 6: AUROC scores against white-box attacks under $\epsilon=4/255$ on CIFAR-10.
Attack BIM PGD DIM MIM AA
AUROC 0.9180 0.9168 0.9062 0.9031 0.9019
Table 7: Impact of data size on distribution estimation: AUROC score against PGD under $\epsilon=4/255$ on CIFAR-10.
Data size 100 250 500 750 1000 2000
AUROC 0.8278 0.9368 0.9866 0.9921 0.9965 0.9948
5.5.3 Detecting White-box Attack
To verify the robustness of our approach to white-box attacks, i.e., both the victim model and the detector are available to the attacker, we conduct the experiment on CIFAR-10. Following [3], a modified objective of attacks is:
$\max_{\boldsymbol{\eta}}\mathcal{L}(f(\hat{\boldsymbol{\mathrm{x}}}),y)+\alpha\cdot\mathcal{L}(f_{d}(\hat{\boldsymbol{\mathrm{x}}}),y_{d}),\ s.t.||\boldsymbol{\eta}||\leq\epsilon$ (12)
where $y$ is the ground-truth label, $y_{d}$ = 0 indicates that $\hat{\boldsymbol{\mathrm{x}}}$ is an adversarial example, and $\alpha=1$ is a constant balancing between the attack of the classifier $f$ and detector $f_{d}$. As shown in Table 6, our method achieves a high average AUROC of 0.9092, demonstrating its robustness against white-box attacks. This may be attributed to the comprehensive open covering formed by perturbation forgery, which effectively encompasses adversarial data from white-box attacks, making it harder to mislead the trained detector.
Figure 4: Impact of the initial attacks (AUROC) on CIFAR-10 under $\epsilon=4/255$. Y-axis: initial attack. X-axis: testing attack.
5.5.4 Impact of the Initial Attack
Our proposed perturbation forgery requires a commonly used attack to construct the open covering. To assess the impact of the initial attack choice, we train detectors using distributions estimated from various initial attacks and test them against other attacks on CIFAR-10. As shown in Figure 4, the detectors consistently achieve a high AUROC score above 0.99, indicating that our method does not depend on a specific attack.
5.5.5 Data size for Distribution Estmation
To assess the influence of data size for distribution estimation, we train detectors using distributions estimated from various data sizes and test them against PGD on CIFAR-10. As shown in Table 7, our method only requires a small amount of data ($\geq 500$ samples) to estimate the adversarial noise distribution so as to train an effective detector. Notably, using fewer than 500 samples leads to a significant decrease in detection performance.
6 Conclusion
@@
In this paper, we explore and define the proximity relationship between adversarial noise distributions by modeling these distributions as multivariate Gaussian. Based on our assumption and definition, we establish a perturbation method to craft proximal distributions from a given distribution and demonstrate the existence of an open covering of adversarial noise distributions. After that, we propose the core idea of obtaining a detector with strong generalization abilities against various types of adversarial attacks by training on the open covering of adversarial noise distributions. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, agnostic to any specific models. Extensive experiments confirm the effectiveness of our approach and demonstrate its strong generalization ability against unseen adversarial attacks, with a satisfactory inference time overhead.